Help! I am stuck on an equation. I need to factor:
4y squared + 7y - 2
I amd just really stuck with how to do this. Even anyone could show me the proper way to factor I would be very grateful.
Thank you.

Answer by checkley71(8405)

4y^2+7y-2=0 or (4y-1)(y+2)=0 or 4y-1=0 or 4y=1 or y=1/4 & y+2=0 or y=-2